# Thermodynamics of Kerr black hole: Tsallis-Cirto composition law and entropy quantization

### v3

@@ Abstract
The processes of splitting and merging of black holes obey the composition law generated by the Tsallis-Cirto $\delta=2$ statistics. The same composition law expresses the full entropy of the Reissner-Nordström black hole via the entropies of its outer and inner horizons. He we apply this composition law to the entropy of the Kerr black hole. As distinct from Reissner-Nordström black hole, where the full entropy depends only on mass $M$ and does not depend on its charge $Q$, the entropy of Kerr black hole is the sum of contributions from its mass $M$ and angular momentum $J$, i.e. $S(M,J)=S(M,0)+4\pi\sqrt{J(J+1)}$. Here $S(M,0)$ is the entropy of the Schwarzschild black hole. This demonstrates that when the Kerr black hole with $J\gg 1$ absorbs or emits a massless particle with spin $s_{z}=\pm 1/2$, its entropy changes by $|\Delta S|=2\pi$. We also considered the quantization of entropy suggested by the toy model, in which the black hole thermodynamics is represented by the ensemble of the Planck-scale black holes – Planckons.
@@
I Introduction. Composition rule and Tsallis-Cirto $\delta=2$ entropy
The black hole entropy $S_{\rm BH}(M)=4\pi GM^{2}$ is non-extensive with the special type of composition. In the process of the splitting of the black hole with mass $M$ into two smaller black holes with masses $M_{1}$ and $M_{2}$, with $M_{1}+M_{2}=M$, the entropy obeys the following rule:
@@
It was shownVolovik2025TC that this composition law can be extended to the white hole horizons with negative entropy and to the multi-horizon systems with positive and negative entropies. For example, if the processes of splitting and merging include both the black holes with positive entropy and the white holes with negative entropy, the Eq.(1) becomes
@@
II Composition rule for Reissner-Nordström black hole
Another example is the composition laws for Reissner-Nordström (RN) black hole, which expresses the full entropy of the RN black holes in terms of the entropies of its outer and inner horizons:Volovik2025TC
@@ II Composition rule for Reissner-Nordström black hole
In our earlier paper,Volovik2021 it was suggested that the similar situation may happen for the rotating black hole, where the adiabatic change of the parameter $\hbar$ was considered and it was concluded that the entropy of the Kerr black hole is only determined by its mass. However, this is not so, see next Section. The reason why this conclusion is incorrect is that changing the parameter $\hbar$ at a fixed angular momentum $L=\hbar J$ is not an adiabatic process, since in this process the quantum number $J$ experiences discrete quantum jumps.
III Entropy quantization from composition rule for Kerr black hole
@@
$S_{+}=2\pi GM^{2}\left(1+\sqrt{1-J^{2}/M^{4}G^{2}}\right)^{2}\,$, (10)
@@
$|S_{-}|=2\pi GM^{2}\left(1-\sqrt{1-J^{2}/M^{4}G^{2}}\right)^{2}\,.$ (11)
@@
IV Entropy quantization in Planckon toy model
@@
Here the dimensionless parameter $C$ depends on the choice of the Planckon mass quantum. In Ref. Volovik2025P the mass of Planckon was chosen as the reduced Planck mass, $m_{\rm P}=1/\sqrt{8\pi G}$, which corresponds to $C=1$. If we choose the Planck mass $M_{\rm P}=1/\sqrt{G}$ as the mass quantum, i.e. $M=NM_{\rm P}$, one has $C=8\pi$ and for the Kerr black hole one obtains:
@@
In this toy model the entropy depends on two quantum numbers, $N$ and $J$, with $J=N^{2}$ for the extremal RN black hole in the semiclassical limit $N\gg 1$. However, as distinct from quantization in terms of angular momentum, which is physical, the quantization in terms of Planckon number $N$ is the property of the toy model. The factor $4\pi$ takes place only for the Planckon mass quantum $M_{\rm P}=1/\sqrt{G}$. For this mass quantum, the extreme regime is expressed entirely in integers, $J=N^{2}$, which is quite intriguing. However, for the extremal RN black hole one has $Q=N/\sqrt{\alpha}$, and $N$ cannot be integer due to the presence of the running coupling $\alpha$ in Eq.(6).
Note that this toy model produces the results, which are similar to that of the $SU(N)$ matrix model discussed in Refs.Chu2025 ; Chu2025b , where the horizon of Schwarzschild black hole is identified with the fuzzy sphere and $N$ is the dimension of the Hilbert space of quantum gravity.
V Conclusion
@@
This is different from the thermodynamics of the Reissner-Nordström black hole, where the application of the Tsallis-Cirto $\delta=2$ statistics shows that the full entropy of the RN black hole depends only on mass $M$ and does not depend on its charge $Q$, i.e. $S(M,Q)=S(M,0)$. It will be interesting to apply this approach to the more complicated black holes, such as Kerr-Newman black hole.

### v5

@@ Abstract
The processes of splitting and merging of black holes obey the composition law generated by the Tsallis-Cirto $\delta=2$ statistics. The same composition law expresses the full entropy of the Reissner-Nordström black hole via the entropies of its outer and inner horizons. He we apply this composition law to the entropy of the Kerr black hole. As distinct from Reissner-Nordström black hole, where the full entropy depends only on mass $M$ and does not depend on its charge $Q$, the entropy of Kerr black hole is the sum of contributions from its mass $M$ and angular momentum $J$, i.e. $S(M,J)=S(M,0)+4\pi\sqrt{J(J+1)}$. Here $S(M,0)$ is the entropy of the Schwarzschild black hole. This demonstrates that when the Kerr black hole with $J\gg 1$ absorbs or emits a massless particle with spin $s_{z}=\pm 1/2$, its entropy changes by $|\Delta S|=2\pi$. We also considered the quantization of entropy suggested by the toy model, in which the black hole thermodynamics is represented by the ensemble of the Planck-scale black holes – Planckons. The Tsallis-Cirto composition law is also extended to the entropy of Kerr-Newman black hole.
@@
I Introduction
Hawking radiation from and Bekenstein entropy of the black holes with multiple horizons and the corresponding thermodynamics can be different from that which is obtained by consideration of only the effects of the outer horizon, see e.g. the recent paperSingha2025 and references therein. For the spherically symmetric Reissner-Nordström (RN) black hole with two horizons, several different approaches have been used.Volovik2021 ; Volovik2025TC This includes: (i) coherent Hawking radiations from two horizonz; (ii) macroscopic quantum tunneling; (iii) adiabatic transformations; (iv) singular coordinate transformations; and also (v) the Tsallis-Cirto non-extensive statistics.TsallisCirto2013 ; Tsallis2020 All these approaches demonstrate that the total entropy of the black hole is not determined by the area of the outer horizon. Due to the coherent correlations between the outer and inner horizons, the total entropy of the RN black hole and the temperature of Hawking radiation depend only on mass of the black hole and do not depend on the black hole charge.
In the paper Volovik2021 , an attempt was made to consider the entropy of rotating black holes – the Kerr black hole and the Kerr-Newman black hole. The method of adiabatic transformations was used for this purpose. However, it turned out that this method is inapplicable for the rotating case. Here, we consider the entropy of rotating black holes using non-extensive Tsallis-Cirto statistics. In Sec. 2 we recall the Tsallis-Cirto type statistics, which provides a composition rule for black hole merger and splitting processes, as well as for the combination of the entropies of the inner and outer horizons of an RN black hole. This composition law is applied to Kerr and Kerr-Newman black holes in Sec. III.
II Composition rule and Tsallis-Cirto $\delta=2$ entropy
II.1 Composition rule for merging and splitting of black holes
The black hole entropy $S_{\rm BH}(M)=4\pi GM^{2}$ is non-extensive with the special type of composition. One example is the process of the splitting of the black hole with mass $M$ into two smaller black holes with masses $M_{1}$ and $M_{2}$. For $M_{1}+M_{2}=M$, the entropy obeys the following composition rule:
@@ II.1 Composition rule for merging and splitting of black holes
It was shownVolovik2025TC that this composition law can be extended to the white hole horizons with negative entropy. For example, if the processes of splitting and merging include both the black holes with positive entropy and the white holes with negative entropy, the Eq.(1) becomes
@@
II.2 Composition rule for inner and outer horizons of Reissner-Nordström black hole
Another example of the application of the Tsallis-Cirto non-extensive statistics is provided by the multi-horizon black holes with positive and negative entropies of horizons. For the Reissner-Nordström (RN) black hole, the composition law expresses the full entropy of the RN black holes in terms of the entropies of its outer and inner horizons:Volovik2025TC
@@ II.2 Composition rule for inner and outer horizons of Reissner-Nordström black hole
In our earlier paper,Volovik2021 it was suggested that the similar situation may happen for the rotating black hole, where the adiabatic change of the parameter $\hbar$ was considered and it was concluded that the entropy of the Kerr black hole is only determined by its mass. However, this is not so, see the next Section III. The reason why this conclusion is incorrect is that changing the parameter $\hbar$ at a fixed angular momentum $L=\hbar J$ is not an adiabatic process, since in this process the quantum number $J$ experiences discrete quantum jumps.
III Rotating black holes
III.1 Entropy of Kerr black hole from composition law
@@
$S_{+}=2\pi GM^{2}\left(1+\sqrt{1-J^{2}/M^{4}G^{2}}\right)\,$, (10)
@@
$|S_{-}|=2\pi GM^{2}\left(1-\sqrt{1-J^{2}/M^{4}G^{2}}\right)\,.$ (11)
@@ III.2 Entropy quantization for Kerr black hole
III.3 Entropy of Kerr-Newman black hole from the composition law
Let us apply the composition rule in Eqs. (5) and (9) to the more complicated black hole – the Kerr-Newman black hole with angular momentum $J$ and charge $Q$. From the entropies of inner and outer horizons of the Kerr-Newman black hole one obtains:
$S(M,J,Q)=S(M,0,0)+2\pi\left(\sqrt{\alpha^{2}Q^{4}+4J^{2}}-\alpha Q^{2}\right).$ (17)
For $Q=0$ this transforms into the equation (12) for the entropy of the Kerr black hole, and for $J=0$ this transforms into the equation (8) for the entropy of the RN black hole.
The entropy of Kerr-Newman black hole depends on the running coupling $\alpha$, but this dependence disappears in the RN limit, i.e. at $J=0$, where the entropy depends only on mass $M$. The reason is that at $J=0$ we can adiabatically transform the RN black hole by slow change of this parameter $\alpha$ to zero at fixed mass $M$ and fixed quantum number $Q$. Since in the adiabatic process the entropy of the RN black hole does not change, it is the same as for the Schwarzschild black hole with the same mass $M$.Volovik2021 At $J\neq 0$ such an adiabatic process is absent, since there are too many parameters that need to be fixed.
III.4 Entropy quantization in Planckon toy model
@@
Here the dimensionless parameter $C$ depends on the choice of the Planckon mass quantum. In Ref. Volovik2025P the mass of Planckon was chosen as the reduced Planck mass, $m_{\rm P}=1/\sqrt{8\pi G}$, which corresponds to $C=1$. If we choose the Planck mass $M_{\rm P}=1/\sqrt{G}$ as the mass quantum, i.e. $M=NM_{\rm P}$, one has $C=8\pi$.
All this suggests that for the Kerr black hole and with $M=NM_{\rm P}$ one would have:
@@
In this toy model the entropy depends on two quantum numbers, $N$ and $J$, with $J=N^{2}$ for the extremal RN black hole in the semiclassical limit $N\gg 1$. Unlike quantization by angular momentum $J$, quantization by Planckon number $N$ is the property of a toy model. The factor $4\pi$ takes place only for the Planckon mass quantum $M_{\rm P}=1/\sqrt{G}$. Nevertheless, for this mass quantum, the extreme regime is expressed entirely in integers, $J=N^{2}$, which is quite intriguing. Note that this toy model produces the results, which are similar to that of the $SU(N)$ matrix model discussed in Refs.Chu2025 ; Chu2025b , where the horizon of Schwarzschild black hole is identified with the fuzzy sphere and $N$ is the dimension of the Hilbert space of quantum gravity.
For the charge black holes, the quantization is violated by the parameter $\alpha$ – the running coupling. For example, the running coupling $\alpha$ in Eq.(6) enters the condition for extremality, $N=\sqrt{\alpha}Q$, and thus NN cannot be integer. The running coupling $\alpha$ also enters the entropy of the Kerr-Newman black hole in Eq.(17).
IV Conclusion
@@
This is different from the thermodynamics of the Reissner-Nordström black hole, where the application of the Tsallis-Cirto $\delta=2$ statistics shows that the full entropy of the RN black hole depends only on mass $M$ and does not depend on its charge $Q$, i.e. $S(M,Q)=S(M,0)$. This agrees with possibility to adiabatically transform the entropy of the RN black hole to the entropy of the Schwarzschild black hole by adiabatically reducing the fine structure constant $\alpha$ to zero.
Using the Tsallis-Cirto $\delta=2$ statistics we also obtained the entropy of the Kerr-Newman black hole with its inner and outer horizons, $S(M,J,Q)=S(M,0,0)+2\pi\left(\sqrt{\alpha^{2}Q^{4}+4J^{2}}-\alpha Q^{2}\right)$. As distinct from the RN black hole with $J=0$, this entropy does depend on $\alpha$ parameter. This demonstrates that while there is an adiabatic path between the Reissner-Nordström and Schwarzschild black holes, there is no adiabatic path between the Kerr-Newman and Kerr black holes.